OpenStudy (anonymous):

Given field is in the shape of a triangle with side lengths 19 meters, 26 meters and 36 meters. Find the largest angle (measured in degrees) in any of the three corners of the field. Help me pls

OpenStudy (jdoe0001):

using the Law of Cosines \(\bf a^2 = b^2+c^2-2bc(cos(a)) \implies \cfrac{a^2-b^2-c^2}{-2bc} = cos(a)\)

OpenStudy (anonymous):

yea thats what i used too

OpenStudy (jdoe0001):

say for example, let's do the side opposite to the 36m in length, because that one with the longer side opposite to it, will be the obtuse angle

OpenStudy (anonymous):

yea

OpenStudy (jdoe0001):

so you got it then :)

OpenStudy (anonymous):

i did the whole thing and kept getting a negative number, but ill try again

OpenStudy (anonymous):

i would help if you could go thru the whole thing just to be sure...

OpenStudy (jdoe0001):

keep in mind that \(\bf -1 \le cosine \le 1\) so negative is fine, so long is \(\bf \le -1\)

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\bf \cfrac{36^2-19^2-26^2}{-2(19)(26)} = cos(a)\)

OpenStudy (anonymous):

i could have put the bottom part in any order right?

OpenStudy (jdoe0001):

yes, they're just factor, so yes

OpenStudy (jdoe0001):

\( \bf \cfrac{36^2-19^2-26^2}{-2(19)(26)} = cos(a)\\ cos^{-1}(\text{some value}) = cos^{-1}(cos(a))\)

OpenStudy (anonymous):

i think i forgat the negative along the way dats why

OpenStudy (anonymous):